OpenStudy (anonymous):

what is the equation of the line that passes through the points (-1, 3/2) and (1/2 , 1/2) .

OpenStudy (nenadmatematika):

y-3/2=((1/2-3/2)/(1/2+1))(x+1) y-3/2=-2/3(x+1) y=(-2/3)x+5/6